If f(x)= 1/2x+6 find the equation of the line for the vertical asymptote and the horizontal asymptote.
Answer by lwsshak3(11628)

If f(x)= 1/2x+6 find the equation of the line for the vertical asymptote and the horizontal asymptote
**
f(x)= 1/(2x+6)
To find the vertical asymptote, set the denominator=0, then solve for x.
2x+6=0
2x=-6
x=-3 (Equation of vertical asymptote)
..
If the degree of the numerator is less than the degree of the denominator, as in this case, the horizontal asymptote is the x-axis or y=0. Just by inspection, you should see that as x increases to infinity, f(x) becomes smaller and approaches zero, the x-axis, but never reaches it.